OpenStudy (anonymous):

Check my answer. Solve 44 = 14 - 2a. My answer: first we subtract 14 from both sides, 44 - 14 = 14 - 14 - 2a = 30 = - 2a, now we divide both sides by - 2. 30/ - 2 = - 2a/ - 2 = - 15 = a.

OpenStudy (anonymous):

That's correct. You can check yourself by pluging in your answer instead of a: \[14 - 2a = 44\] \[14 - 2(-15) = 44\] \[14 + 30 = 44\] \[44 = 44\]

OpenStudy (anonymous):

Thanks